Math 141H, Section 1

Honors Calculus II

Stephen G. Simpson 333 McAllister, 863-0775

                     simpson@math.psu.edu

January 8, 2002

My office hours this semester are Monday Wednesday Friday 1:55–2:50 PM, and by appointment.

Homework will be assigned as in the regular Math 141 syllabus and collected on Wednesdays. No late homework will be accepted. Only a small number of homework problems will be graded. These will be chosen from among the more advanced problems.

Three sets of Bonus Problems will be assigned.

There will be no quizzes or evening exams. We will have four in-class Midterm Exams: January 23, February 20, March 20, April 10 (all Wednesdays). Makeup exams will be given during office hours a week later, and only in case of a valid, documented excuse, such as illness. Personal business is not a valid excuse.

There will be a Final Exam, to be scheduled during Final Exam Week, April 29 to May 3. Please do not make arrangements to leave town before the end of Final Exam Week.

Grades will be based on a total of 500 points: 50 for each Midterm Exam, 150 for the Final Exam, 100 for homework, 50 for Bonus Problems.

Some useful web sites: Undergraduate Mathematics at Penn State:

http://www.math.psu.edu/UG/.

Math 141:

http://www.math.psu.edu/m141/.

Math 141H Section 1 (this section):

http://www.math.psu.edu/simpson/courses/math141/.